A rectangle has a perimeter of 100 feet. Express its area in terms of one of its sides, x
Answer by ankor@dixie-net.com(22740)

A rectangle has a perimeter of 100 feet.
Express its area in terms of one of its sides, x
:
Let L = the length
Let x = the width
:
The perimeter:
2L + 2x = 100
simplify, divide by 2
L + x = 50
L = (50-x)
:
The area:
A = L * x
Replace L with (50-x)
A = x(50-x)
A = -x^2 + 50x